OpenStudy (ayeeeitskateee):

If the density is defined as mass per volume, what is the volume of 83.0g of ether if the density of either is 0.70 g/mL? Show your work and explain to me how to do this. Please don't just give me an answer, I need an explanation for future reference. (I will most likely write it in my notebook) I will medal and fan whoever helps the most, thanks!!

OpenStudy (ayeeeitskateee):

@texaschic101 @whpalmer4 Do you guys know how to do this? If not, please tag somebody who does. Thank you.

OpenStudy (ayeeeitskateee):

@nincompoop

OpenStudy (anonymous):

\[density=\frac{ mass }{ volume }\] aka \[\rho=\frac{m}{V}\] The mass and density are given, so substitute them into the equation to solve for the volume. \[0.70~g/mL=\frac{ 83.0~g }{ V }\]

OpenStudy (ayeeeitskateee):

Would I have to multiply 83g on each side to cancel it out and get the V by itself?

OpenStudy (anonymous):

dividebecause 83 is in the numerator, but it might be easier to cross multiply.